OpenStudy (anonymous):

What is the n coordinate of the solution? 2m - 3n = 6 and m = 3 - 2n

OpenStudy (texaschic101):

m = 3 - 2n now sub that in for m in the other equation, then solve for n. Let me get you started... 2m - 3n = 6 2(3 - 2n) - 3n = 6 can you take it from here ?

OpenStudy (anonymous):

yea thank you

OpenStudy (texaschic101):

happy to help :)